Presented below are condensed financial statements adapted from those of two actual companies competing in the pharmaceutical industry—Johnson and Johnson (J&J) and Pfizer, Inc. ($ in millions, except per share amounts).

Balance Sheets
($ in millions, except per share data)
J&J Pfizer
Assets:
Cash $ 16,779 $ 11,244
Short-term investments 6,144 12,400
Accounts receivable (net) 8,894 11,095
Inventories 5,744 9,603
Other current assets 5,790 5,655
Current assets 43,351 49,997
Property, plant, and equipment (net) 15,542 23,983
Intangibles and other assets 19,518 72,843
Total assets $ 78,411 $ 146,823
Liabilities and Shareholders' Equity:
Accounts payable $ 7,126 $ 4,761
Short-term notes 4,835 12,514
Other current liabilities 9,343 14,238
Current liabilities 21,304 31,513
Long-term debt 4,955 7,755
Other long-term liabilities 6,959 23,954
Total liabilities 33,218 63,222
Capital stock (par and additional paid-in capital) 5,520 69,450
Retained earnings 48,683 47,062
Accumulated other comprehensive income (loss) (910 ) 275
Less: Treasury stock and other equity adjustments (8,100 ) (33,186 )
Total shareholders' equity 45,193 83,601
Total liabilities and shareholders' equity $ 78,411 $ 146,823
Income Statements
Net sales $ 50,422 $ 53,748
Cost of goods sold 14,192 11,848
Gross profit 36,230 41,900
Operating expenses 21,715 30,438
Other (income) expense—net (625 ) 3,850
Income before taxes 15,140 7,612
Tax expense 4,542 2,284
Net income $ 10,598 $ 5,328 *
Basic net income per share $ 3.22 $ 0.38

* This is before income from discontinued operations.

Evaluate and compare the two companies by responding to the following questions.

Note: Because two-year comparative statements are not provided, you should use year-end balances in place of average balances as appropriate.

Required:
1. For both companies, compute the ratios below.

For both companies, compute the ratios below.

J&J Pfizer
Receivables Turnover (#.##) times times
Average Collection Period (#) days days
Inventory Turnover (#.##) times times
Average Days in Inventory (#) days days
Profit Margin (#.##)% % %
Asset Turnover (#.###) times times
Return on Assets (#.#)% % %
Equity Multiplier (#.##)
Return on Shareholders' Equity (#.#)% % %

2. Evaluate and compare the two companies.
  

Analysis
Which of the two companies appears more efficient in collecting its accounts receivable and managing its inventory?
Which of the two firms had greater earnings relative to resources available?
Have the two companies achieved their respective rates of return on assets with similar combinations of profit margin and turnover?
From the perspective of a common shareholder, which of the two firms provided a greater rate of return?
From the perspective of a common shareholder, which of the two firms appears to be using leverage more effectively to provide a return to shareholders above the rate of return on assets?
Answer #1

Receivables Turnover = Sale / Accounts Receivable

J&J = 50,422 / 8,894 = 5.67

Pfizer = 53,748 / 11095 = 4.84

Average Collection Period = 365 / Receivables Turnover

J&J = 365 / 5.67 = 64.37 Days

Pfizer = 365 / 4.84 = 75.41 Days

J&J appears more efficient in collecting its accounts receivable

Inventory Turnover = Cost of goods sold / Inventory

J&J = 14192 / 5744 = 2.47

Pfizer = 11848 / 9603 = 1.233

Average Days in Inventory = 365 / Inventory Turnover

J&J = 365 / 2.47 = 147.77 Days

Pfizer = 365 / 1.233 = 296.03 Days

J&J appears more efficient in managing its inventory

Profit Margin = Net Income / Net sales

J&J = 10598 / 50422 = 21.02%

Pfizer = 5328 / 53748 = 9.91%

J&J had greater earnings relative to resources available

Asset Turnover = Sales / Total Assets

J&J = 50422 / 78411 = 0.643

Pfizer = 53748 / 146823 = 0.366

Return on Assets = Net Income / Total Assets

J&J = 10598 / 78411 = 13.52%

Pfizer = 5328 / 146823 = 3.63%

No the two companies have not achieved their respective rates of return on assets

Equity Multiplier = Total Assets / Total Equity

J&J = 78411 / 45193 = 1.74

Pfizer = 146823 / 83601 = 1.76

Return on Shareholders' Equity = Net Income / Total Equity

J&J = 10598 / 45193 = 23.45%

Pfizer = 5328 / 83601 = 6.37%

From the perspective of a common shareholder J&J provided a greater rate of return